OpenStudy (anonymous):

Michelle rents a movie for a flat fee of $1.50 plus an additional $1.25 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Michelle has the movie.

OpenStudy (anonymous):

same one right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

c(x) = 1.50 + 1.25x c(x) = 1.50x + 1.25 c(x) = 2.75 c(x) = (1.50 + 1.25)x those are the answer choices

OpenStudy (anonymous):

if she keeps it for one night she pays $1.50 + $1.25 = $3.75 if she keeps it for two nights she pays $1.50 + $1.25 + $1.25 = $4.00

OpenStudy (anonymous):

ah i see if she keeps it for \(x\) nights she pays \(1.50 + 1.25x\) dollars

OpenStudy (anonymous):

so the first one?

OpenStudy (anonymous):

The first one is correct:)

OpenStudy (anonymous):

thank you